OpenStudy (anonymous):

Rationalise and simplify 4 )5 divided by )3 ... this ) is square root. How do you work it out?

OpenStudy (neer2890):

\[\frac{ 4\sqrt{5} }{ \sqrt{3} }\] to rationalise the denominator, it should not contain any square roots because square root of a number is not always rational. so to rationalise the denominator multiply it by )3 \[\frac{ 4\sqrt{5} }{ \sqrt{3} }*\frac{ \sqrt{3} }{ \sqrt{3} }=\frac{ 4\sqrt{15} }{ 3 }\]

OpenStudy (anonymous):

can you simplfy the 15 and the 3 so the equation ends up 4√5? @neer2890

OpenStudy (neer2890):

how can you say that \[\frac{ 4\sqrt{5} }{ \sqrt{3} }=4\sqrt{5}\] keep in mind , rationalizing does not changes the actual values.

OpenStudy (yanasidlinskiy):

By the way, @Mmonika \(\Huge\bf \color{yellow}{Welcome~to~OpenStudy!!}\hspace{-310pt}\color{cyan}{Welcome~to~OpenStudy!!}\hspace{-307.1pt}\color{midnightblue}{Welcome~to~\color{purple}{Open}}\color{blue}{Study!!}\)